The measure of an angle's supplement is 24 less than twice the measure of the angle. Find the measure of the angle and its supplement.
I've tried:
setting up an equation but im not sure if its correct...
2x-24=180
x=102

Answer by checkley71(8403)

x+(2x-24)=180
3x-24=180
3x=180+24
3x=204
x=204/3
x=68 degrees for the smaller angle
180-68=112 degrees for the larger angle.
proof
68+(2*68-24)=180
68+136-24=180
204-24=180
180=180